40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 22 23:16:54 2020, Koji, Update, General, HEPA Inspection 11x
    Reply  Tue Sep 22 23:17:51 2020, Koji, Update, General, HEPA Inspection 6x
       Reply  Wed Sep 30 17:12:24 2020, gautam, Update, General, HEPA blower startup capacitor replacement 
          Reply  Wed Sep 30 19:45:56 2020, rana, Update, General, HEPA blower startup capacitor replacement 
Message ID: 15604     Entry time: Wed Sep 30 17:12:24 2020     In reply to: 15598     Reply to this: 15605
Author: gautam 
Type: Update 
Category: General 
Subject: HEPA blower startup capacitor replacement 

[JV, GV]

The HEPAs work again. After running the HEPAs for ~1 hour, I checked the particle count on the PSL table - the meter registered 0 for both 0.3 um and 0.5 um. So I decided to turn the NPRO back on, at ~1730 local time. The PMC and IMC were readily locke, so the basic interferometer functionality is returned, and we can now go ahead with (i) vent prep (ii) air BHD tests and (iii) IMC debuggin as was discussed on the call today. The earth is shaking, but nothing serious so far, I will resume alignment of the interferometer later in the evening when hopefully things have calmed down a bit more...

Procedure:

  1. Turned off mains switch on the NW corner of the PSL enclosure. Then, disconnected the power cables from mains to Variac and from Variac to HEPAs. Made sure both HEPAs were set to "OFF".
  2. With confidence that no AC power was reaching the motors, I removed the pre-filters, and removed the old startup capacitors. These don't have a polarity, but I marked the cable that was connected to the left terminal of the capacitor when the cap is viewed with the label facing you, in the interest of changing as few things as possible.
  3. The two new capacitors were measured with the LCR meter - the meter registered ~7.5 uF, as expected. Unsurprisingly, the old capacitors that were removed didn't register any reading on the LCR meter. The terminals weren't shorted, but I don't know what the failure mode for this kind of capacitor is.
  4. The two new capacitors were installed. Then, I tested the system by undoing all the changes in bullet #1. We found that the Variac needs to be set to 100% for the motors to startup.
  5. The motor speed was found to vary as the Variac dial was turned. FWIW, at the "nominal" setting of 33% on the Variac (when we run the interferometer), I could see both fan blades were turning, but the flow was low enough that you couldn't hear any wind (at least, neither Jordan nor I could).
  6. Turned off the mains agian, and cleaned up - restored the insulating rubber sleeve on the capacitor leads, and re-installed the pre-filters on the HEPA blowers. Then we turned both blowers back on. 

Note that the many other issues Koji noted in the preceeding elog (e.g. flaky wiring) have not been addressed.

Flow measurements:

Chub kindly provided us with an electronic anemometer. With the meter held directly against the HEPA filter inside the enclosure, we measured ~700 cfm of airflow on each of the two HEPAs, with the Variac set to 100% and the HEPAs themselves set to "High". With the Variac at 50%, the flow drops to ~160 cfm. At the nominal setting of 33%, the meter didn't register any flow. I don't know what the spec'd flow rate is for this combination of blower + filter, but Jordan says similar units in Downs register ~1500 cfm at the "High" setting. The two protable (similarly sized) HEPA units in the 40m, one at ITMY and one at ETMY, register ~900 cfm and ~1100 cfm respectively, when set to high. So we may want to revisit what the "nominal" HEPA setting should be, in case the filters have become clogged over time. 

Some photos of the HEPA blowers with the pre-filters off and the capacitors switched out may be found here.

ELOG V3.1.3-